What is the ARR, NPV, IRR, and profitability index?Splash Planet is considering purchasing a water park in Atlanta, Georgia, for $1,820,000. The new facility will generate annuiReference Present Value of $1 4 %5%| 6%|7% 8%| 9% | 10%| 12%|14%| 15% Reference Future Value of $1 1% 23% 4%5%6% 7%8%9%10% 1
Answer #1
payback period = amount invested/expected annual return
1820000/460000 = 4 years
ARR = Average annual operating income/average amount invested
460000/1820000 = 25%
net present value = present value of cash flow-investment
present value of cash inflow = 460000*PVIFA(12%, 8YEARS)
present value of cash inflow = 460000*4.968 = 2285280
Net present value = 2285280-1820000 = 465280
calculation of IRR
Years cash flow
0 -1820000
1 460000
2 460000
3 460000
4 460000
5 460000
6 460000
7 460000
8 460000
IRR = 19%
Profitability index = present value of cash inflow/initial investment
Profitability index = 2285280/1820000 = 1.26